OpenStudy (anonymous):

What is the output value for the following function if the input value is 5? y = 4x - 3 42 23 17 2

OpenStudy (anonymous):

17

jimthompson5910 (jim_thompson5910):

Replace x with 5, then evaluate y = 4x-3 y = 4(5) - 3

Parth (parthkohli):

x is the input. y is the output.

Parth (parthkohli):

y = 4(5) - 3 y = ?